# Data-Efficient RLVR via Off-Policy Influence Guidance

### v1

@@ Abstract
Data selection is a critical aspect of Reinforcement Learning with Verifiable Rewards (RLVR) for enhancing the reasoning capabilities of large language models (LLMs). Current data selection methods are largely heuristic-based, lacking theoretical guarantees and generalizability. This work proposes a theoretically-grounded approach using influence functions to estimate the contribution of each data point to the learning objective. To overcome the prohibitive computational cost of policy rollouts required for online influence estimation, we introduce an off-policy influence estimation method that efficiently approximates data influence using pre-collected offline trajectories. Furthermore, to manage the high-dimensional gradients of LLMs, we employ sparse random projection to reduce dimensionality and improve storage and computation efficiency. Leveraging these techniques, we develop Curriculum RL with Off-Policy Influence guidance (CROPI), a multi-stage RL framework that iteratively selects the most influential data for the current policy. Experiments on models up to 7B parameters demonstrate that CROPI significantly accelerates training. On a 1.5B model, it achieves a 2.66x step-level acceleration while using only 10% of the data per stage compared to full-dataset training. Our results highlight the substantial potential of influence-based data selection for efficient RLVR.
@@ 2 Preliminaries
where $\rho_{k,t}^{\pi_{\theta}}=\frac{\pi_{\theta}(x_{k,t}|s_{k,t})}{\pi_{\theta_{\text{old}}}(x_{k,t}|s_{k,t})}$ and $\widehat{A}_{k,t}=\frac{R(\tau_{k})-\widehat{\mathbb{E}}_{\pi_{\theta_{\text{old}}}}[R(\tau)]}{\widehat{\sigma}_{\pi_{\theta_{\text{old}}}}[R(\tau)]}$. $\widehat{\mathbb{E}}$ and $\widehat{\sigma}$ is the empirical mean and standard derivation of the return of trajectories $\{\tau_{k}\}_{k=1}^{K}$ sampled from $\pi_{\theta_{\text{old}}}$ given query $s_{0}$.
@@ 3 Influence Estimation in RLVR
Following the first-order influence function formula (Pruthi et al., 2020), in the RLVR context, given a training prompt $s_{0}$ and a test query $s_{0}^{\prime}$, we use the inner product the policy gradients of $s_{0},s_{0}^{\prime}$ to measure the influence of $s_{0}$ to policy’s performance on $s_{0}^{\prime}$:
@@
Rollout Issue. Unlike supervised learning whose training dynamics are predominantly shaped by labeled data and optimization algorithms (Xia et al., 2024; Gu et al., 2024), RL involves online sampling, making its training process more dynamic and less predictable. Consequently, global data selection strategies (Zhao et al., 2025; Wang et al., 2025) can hardly capture the evolving characteristics of policy learning in RLVR. Therefore, we seek to dynamically evaluate the influence of each data point $s_{0}$ conditioned on the current policy $\pi_{\theta}$, thereby enabling effective online utility estimation. However, accurate influence estimation typically requires computing the policy gradient for each $s_{0}$, which demands rollouts over multiple trajectories (Eq. (3)). The substantial computational costs and latency associated with these rollouts present a significant barrier to real-time influence estimation in LLMs.
@@ 3.1 Off-Policy Gradient for Rollout-Free Estimation
The experimental results indicate that the off-policy gradient can approximate the on-policy gradient to a certain extent; please refer to Appendix F for further details.
@@ 3.3 Final Solution
In order to eliminate the bias of the gradient norm caused by different lengths and pass rates, following Xia et al. (2024), we normalize the gradient features before inner product, which is equivalant to computing the cossim.
@@ 4.1 Data Selection with POPI
For multiple validation sets, we need to consider the effectiveness over different validation sets. We use a ranking fusion method Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) which is widely use in IR field to combine the POPI scores over different validation sets. Denote the training set as $\mathcal{D}_{\text{tr}}$, number of $V$ validation sets as $\mathcal{D}_{\text{val},j},j=1,2,...V$, the rank of POPI score of training data $s_{0}$ to $j$-th validation set over the whole training set as $\text{r}_{j}(s_{0})=\text{rank}_{\text{POPI}}(\pi_{\theta};s_{0},\mathcal{D}_{\text{val},j})$, then the fused rank score is defined as:
@@ 5 Experiments
Table 1: Evaluation results of CROPI and other baseline methods across various math datasets. CROPI consistently outperforms existing data selection approaches, achieving state-of-the-art performance on target tasks at different training steps and under various experimental settings. The best results at each training step are highlighted in bold.
@@ 5.1 Setup
We evaluate CROPI on mathematical reasoning tasks using several open-source, instruction-aligned models diversed from different model scales and context sizes: Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct (Qwen et al., 2025), and Deepseek-R1-Distill-Qwen-1.5B (Guo et al., 2025). For simplicity, these models are referred to 1.5B, 7B, and 1.5B-R1, respectively. Our training data consists of 47K unique problems aggregated from GSM8K-Train (Cobbe et al., 2021), MATH-Train(Hendrycks et al., 2021), and DeepScaleR-Preview-Dataset (Luo et al., 2025). We assess performance on a suite of standard benchmarks, including GSM8K-Test, MATH-Test, Gaokao2023EN (MARIO-Math, 2024), OlympiadBench (He et al., 2024), AMC23 (math ai, 2024b) and AIME24 (math ai, 2024a). For influence-based data selection, a small validation set (max 100 examples) is sampled from a subset of these test sets. We refer to tasks in which this validation set is employed for data selection in CROPI as "Targeted," while other test sets are designated as untargeted tasks ("Untar.").
We compared CROPI with several data selection baselines in selection ratio $\alpha=0.1$: Learnability (Bae et al., 2025), Pass Rate (Yu et al., 2025), Influence Function (Pruthi et al., 2020) for global-level data selection, and the DAPO (Yu et al., 2025) for batch-level data selection. Due to computational constraints, baseline comparisons are conducted on the 1.5B model. Full details regarding model versions, dataset construction, and hyperparameters are available in Appendix E.
@@ 5.2 Main Results
As shown in Table 1, CROPI consistently outperforms both full-data training and all baselines across different model scales in targeted tasks, demonstrating superior sample and step efficiency. The improvements are particularly pronounced in the early stages of training. For the 1.5B model, CROPI achieves a remarkable $\bm{2.66\times}$ step-level speedup compared to training on the full dataset on targeted tasks, as illustrated in Figure 3. This acceleration is achieved while using only 10% of the training data in each phase. In contrast, while other data selection baselines show some initial gains, their performance plateaus quickly because they rely on a static estimation of data utility and fail to adapt to the evolving policy. Meanwhile, CROPI also exhibits strong generalization capabilities. We observe significant performance gains on "Untargeted" benchmarks—those not used for creating the validation set—indicating that the data selected by CROPI benefits the model’s overall reasoning ability, not just performance on the targeted tasks.
Figure 3: Training curves on 1.5B setting. CROPI surpasses all other baselines and achieves a significant step-level accelerate ratio $\bm{2.66\times}$ compared to full data training while using only 10% of data during each phase.
As can be seen on Table 2, due to the requirement to compute gradients over the entire dataset (after filtering out samples that are completely correct or incorrect), the data selection time for CROPI remains non-negligible without rollouts. However, even after accounting for the selection time, the slowdown factor is only about 0.81x, and CROPI still achieves a 2.16x speedup overall. Moreover, there remains significant room for optimization in our gradient computation process, such as selecting partial subsets, accelerating parallelization, or training a proxy scorer. Thus, the time speedup of CROPI could be further improved, highlighting its considerable potential.
@@
Table 2: Time costs of data selection and training for CROPI in each phase. We denote the number of prompts to process (gradient computation, projection, cossim) in data selection and optimization steps in training stage in brackets. Time cost is computed in 8-GPUs (NVIDIA H100) machine.
@@ 6 Analysis
This chapter takes in-depth analysis in two key computational components of the CROPI framework: random projection and data selection. We also provide a empirical analysis of off-policy gradient estimation in POPI in Appendix F.
@@ 6.1 Analysis on Sparse Random Projection
As mentioned on Section 3, we use sparse random projection for the projection of full-parameter gradients. Specifically, we randomly select a proportion of dimension of the gradient to perform random projection to avoid computation over the entire high-dimensional gradient. We define this proportion as sparse ratio. However, through our experiments, we surprisingly found that this dropout actually improve the preservation of the inner products between gradient feature after random projection.
@@
Figure 4 presents the experimental results under the 1.5B Setting. We observe that when we directly apply random projection to the full gradient (i.e. sparse ratio equals 1) , the precision@10% of similarity ranking is only around 13%, comparable to random selection. However, when the sparse ratio is 0.1, the precision@10% is significantly higher, reaching nearly 80%. This is a counter-intuitive phenomenon: under sparse random projection, the ranking preservation is actually better with less information. We hypothesize that this may be related to the presence of numerical noise in the gradients: random projection can amplify such noise. Sparsity, while masking some information, filters out much of the numerical noise and achieves a better signal-to-noise ratio around a sparse ratio of 0.1. We include more analysis in Appendix G.
@@ 6.2 Analysis on data selected by CROPI
In Appendix H, we provide a detailed breakdown of data source, knowledge catagirues, diversity of selected data. These results demonstrate that CROPI not only achieves strong performance but also offers a degree of interpretability.
@@ 7 Conclusion
Numerical error in gradient computation. As we utilize float16 format to conduct gradient-related calculation, numerical error can not be ignored and might be increased by random projection. We proposed a randomly dropout operation before random projection as a way to gain better rank preservation. Additional details can be found in Appendix G.
@@ Appendix B Derivation of Off-Policy Gradient
However, for multi-step trajectory generation, the variance of the importance weight $\rho_{\theta}(\tau|s_{0})$ can become extremely large. This high variance arises due to the product of many likelihood ratios over the steps, leading to instability and poor sample efficiency in the gradient estimate. We a more widely used off-policy gradient estimator from (Schulman et al., 2015, 2017) that uses importance sampling in token level:
@@ Appendix C Related Works
C.1 Data Selection for RLVF
Recently, how to automatically select high-quality data for RLVF has become a research topic of increasing interest. According to the time granularity of data selection, existing methods can be roughly divided into three categories: global-level, batch-level and phase-level. The first is the global-level granularity, which selects data based on the initial policy over the entire dataset Zhao et al. (2025); Wang et al. (2025). Although this approach is easy to implement, it cannot capture dynamic checkpoint information and often requires warm-up training. The second granularity is at the batch-level, where data selection is performed within each training batch. This method can better capture the dynamics of model training; however, due to the lack of long-term planning, it often introduces high variance that can lead to unstable training. Representative methods include DAPO (Yu et al., 2025) and ODF (Bae et al., 2025). The third category is a compromise, where data selection is performed every certain number of training steps, which we refer to as phase-level (or curriculum) data selection. This approach strikes a balance between capturing training dynamics and enabling long-term planning, but may still face challenges such as unstable training distributions. An example of this method is Xi et al. (2024).
As for the utiltity estimator, previous works use various ways to measure the utility of the data point, mainly focusing on some heuristics around difficulty and uncertainty of the training query $s_{0}$. In order to select prompts with proper difficulties, some researchers use correctness rate of various trajectories as the signal of the difficulty (Li et al., 2025; Yu et al., 2025; Bae et al., 2025; Sun et al., 2025). As an example, Bae et al. (2025) suggests filtering data points with pass rate around $0.5$ to construct the batch for GRPO training. Zhao et al. (2025) estimate prompt difficulty based on the model’s confidence (likelihood), and select moderately difficult prompts for reinforcement learning (RL) training. Another heuristic involves selecting data according to the uncertainty arising from data perturbation. For instance, Wang et al. (2025) train with prompts whose associated historical trajectories exhibit the highest reward variance.
@@ C.2 Influence Functions for LLM Data Selection
Influence function (Koh and Liang, 2017) is a gradient-based data attribution method derived from variantion analysis of objective function. Influence functions and their variants have been widely applied in the Pre-Training (PT) and Supervised Fine-Tuning (SFT) stages due to their strong theoretical guarantees and empirical effectiveness Logan Engstrom (2024); Wang et al. (2024); Gu et al. (2024); Grosse et al. (2023). Logan Engstrom (2024) proposed the datamodel framework leveraging an efficient influence function estimation Park et al. (2023), which is inspired by influence function, to select pretraining data. LESS (Xia et al., 2024) is the first to introduce first-order influence functions into the post-training of large language models (LLMs), thereby improving the training efficiency of supervised fine-tuning (SFT). Wang et al. (2024) further refine LESS by extending it to the batch-level; however, its application remains limited to the SFT stage. Nonetheless, leveraging influence function theory to guide data selection for RLVF remains a challenging problem. This is mainly due to the fact that, in RL settings, rewards and gradients for the data typically require rollouts to obtain, which is computationally expensive for LLMs. A concurrent work Hu et al. (2025) explores online data selection using influence functions in RLHF. However, their setting differs substantially from RLVF, and the definition of the target function for influence estimation is largely heuristic.

### v2

@@ Abstract
Data selection is a critical aspect of Reinforcement Learning with Verifiable Rewards (RLVR) for enhancing the reasoning capabilities of large language models (LLMs). Current data selection methods are largely heuristic-based, lacking theoretical guarantees and generalizability. This work proposes a theoretically-grounded approach using influence functions to estimate the contribution of each data point to the learning objective. To overcome the prohibitive computational cost of policy rollouts required for online influence estimation, we introduce an off-policy influence estimation method that efficiently approximates data influence using pre-collected offline trajectories. Furthermore, to manage the high-dimensional gradients of LLMs, we employ sparse random projection to reduce dimensionality and improve storage and computation efficiency. Leveraging these techniques, we develop Curriculum RL with Off-Policy Influence guidance (CROPI), a multi-stage RL framework that iteratively selects the most influential data for the current policy. Experiments on models up to 7B parameters demonstrate that CROPI significantly accelerates training. On a 1.5B model, it achieves a $2.66\times$ step-level acceleration while using only 10% of the data per stage compared to full-dataset training. Our results highlight the substantial potential of influence-based data selection for efficient RLVR. Code is available at https://github.com/thu-coai/CROPI.
@@ 2 Preliminaries
where $\rho_{k,t}^{\pi_{\theta}}=\frac{\pi_{\theta}(x_{k,t}|s_{k,t})}{\pi_{\theta_{\text{old}}}(x_{k,t}|s_{k,t})}$ and $\widehat{A}_{k,t}=\frac{R(\tau_{k})-\widehat{\mathbb{E}}_{\pi_{\theta_{\text{old}}}}[R(\tau)]}{\widehat{\sigma}_{\pi_{\theta_{\text{old}}}}[R(\tau)]}$. $\widehat{\mathbb{E}}$ and $\widehat{\sigma}$ denote the empirical mean and standard deviation of the returns of trajectories $\{\tau_{k}\}_{k=1}^{K}$ sampled from $\pi_{\theta_{\text{old}}}$ given query $s_{0}$.
@@ 3 Influence Estimation in RLVR
Following the first-order influence function formula (Pruthi et al., 2020), in the RLVR context, given a training prompt $s_{0}$ and a test query $s_{0}^{\prime}$, we use the inner product of the policy gradients of $s_{0}$ and $s_{0}^{\prime}$ to measure the influence of $s_{0}$ on the policy’s performance on $s_{0}^{\prime}$:
@@
Rollout Issue. Unlike supervised learning whose training dynamics are predominantly shaped by labeled data and optimization algorithms (Xia et al., 2024; Gu et al., 2024), RL involves online sampling, making its training process more dynamic and less predictable. Consequently, global data selection strategies (Zhao et al., 2025; Wang et al., 2025) can hardly capture the evolving characteristics of policy learning in RLVR. Therefore, we seek to dynamically evaluate the influence of each data point $s_{0}$ conditioned on the current policy $\pi_{\theta}$, thereby enabling effective online utility estimation. However, accurate influence estimation typically requires computing the policy gradient for each $s_{0}$, which demands rollouts over multiple trajectories (Eq. (3)). The substantial computational costs and latency associated with these rollouts present a significant barrier to real-time influence estimation in LLMs. In our 1.5B setting with batch size 128 and maximum response length 8192, one rollout step takes 319.80s on average, while one forward+backward step takes only 35.91s, making rollout about 8.91x more expensive. This gap is also consistent with the underlying hardware pattern: forward+backward is mostly compute-bound, while autoregressive decoding is largely memory-bandwidth-bound. Therefore, a rollout-free utility estimator is not merely a convenience but a key requirement for making online data selection practical; we provide additional timing details in Appendix F.
@@ 3.1 Off-Policy Gradient for Rollout-Free Estimation
The experimental results indicate that the off-policy gradient can approximate the on-policy gradient to a certain extent; please refer to Appendix G for further details.
@@ 3.3 Final Solution
In order to eliminate the bias of the gradient norm caused by different lengths and pass rates, following Xia et al. (2024), we normalize the gradient features before taking the inner product, which is equivalent to computing cosine similarity.
@@ 4.1 Data Selection with POPI
For multiple validation sets, we need to consider the effectiveness over different validation sets. We use Reciprocal Rank Fusion (RRF) (Cormack et al., 2009), which is widely used in information retrieval, to combine the POPI scores over different validation sets. Denote the training set as $\mathcal{D}_{\text{tr}}$, number of $V$ validation sets as $\mathcal{D}_{\text{val},j},j=1,2,...V$, the rank of POPI score of training data $s_{0}$ to $j$-th validation set over the whole training set as $\text{r}_{j}(s_{0})=\text{rank}_{\text{POPI}}(\pi_{\theta};s_{0},\mathcal{D}_{\text{val},j})$, then the fused rank score is defined as:
@@ 5 Experiments
Table 1: Evaluation results of CROPI and other baseline methods across various math datasets. CROPI consistently outperforms existing data selection approaches, achieving state-of-the-art performance on target tasks at different training steps and under various experimental settings. The best results at each training step are highlighted in bold. All Acc.@step values are computed from moving-average-smoothed evaluation curves with a window size of $5$.
@@ 5.1 Setup
We evaluate CROPI on mathematical reasoning tasks using several open-source, instruction-aligned models spanning different model scales and context sizes: Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct (Qwen et al., 2025), and Deepseek-R1-Distill-Qwen-1.5B (Guo et al., 2025). For simplicity, these models are referred to 1.5B, 7B, and 1.5B-R1, respectively. Our training data consists of 47K unique problems aggregated from GSM8K-Train (Cobbe et al., 2021), MATH-Train(Hendrycks et al., 2021), and DeepScaleR-Preview-Dataset (Luo et al., 2025). We assess performance on a suite of standard benchmarks, including GSM8K-Test, MATH-Test, Gaokao2023EN (MARIO-Math, 2024), OlympiadBench (He et al., 2024), AMC23 (math-ai, 2024b) and AIME24 (math-ai, 2024a). For influence-based data selection, a small validation set (max 100 examples) is sampled from a subset of these test sets. We refer to tasks in which this validation set is employed for data selection in CROPI as "Targeted," while other test sets are designated as untargeted tasks ("Untar.").
We compared CROPI with several data selection baselines in selection ratio $\alpha=0.1$: Learnability (Bae et al., 2025), Pass Rate (Yu et al., 2025), Influence Function (Pruthi et al., 2020) for global-level data selection, and the DAPO (Yu et al., 2025) for batch-level data selection. Due to computational constraints, baseline comparisons are conducted on the 1.5B model. Full details regarding model versions, dataset construction, and hyperparameters are available in Appendix E. Unless otherwise stated, all reported Acc.@step numbers and training curves use a moving average with a sliding window of size $5$ to reduce the short-term variance of RL training.
@@ 5.2 Main Results
As shown in Table 1, CROPI consistently outperforms both full-data training and all baselines across different model scales in targeted tasks, demonstrating superior sample and step efficiency. The improvements are particularly pronounced in the early stages of training. For the 1.5B model, CROPI achieves a remarkable $\bm{2.66\times}$ step-level speedup compared to training on the full dataset on targeted tasks, as illustrated in Figure 3. Here, using only 10% of the training data in each phase means that at every phase we re-score the full training pool with the current policy, select the top 10% prompts, and run GRPO only on that refreshed subset. In contrast, while other data selection baselines show some initial gains, their performance plateaus quickly because they rely on a static estimation of data utility and fail to adapt to the evolving policy. Meanwhile, CROPI also exhibits strong generalization capabilities. We observe significant performance gains on "Untargeted" benchmarks—those not used for creating the validation set—indicating that the data selected by CROPI benefits the model’s overall reasoning ability, not just performance on the targeted tasks. Additional analyses in Appendix E.4 and Appendix E.5 further show that CROPI reaches its best smoothed accuracy earlier than standard GRPO and that replacing influence-based selection with random phase-wise selection leads to a clear performance drop.
Figure 3: Training curves on the 1.5B setting. Curves are smoothed with a moving average of window size $5$. CROPI surpasses all other baselines and achieves a significant step-level acceleration ratio of $\bm{2.66\times}$ compared to full-data training while using only 10% of the data in each phase.
As can be seen in Table 2, due to the requirement to compute gradients over the entire dataset (after filtering out samples that are completely correct or incorrect), the data selection time for CROPI remains non-negligible even without new rollouts. However, even after accounting for the selection time, the slowdown factor is only about 0.81x, and CROPI still achieves a 2.16x speedup overall. Moreover, there remains significant room for optimization in our gradient computation process, such as selecting partial subsets, accelerating parallelization, or training a proxy scorer. Thus, the time speedup of CROPI could be further improved, highlighting its considerable potential. For the 1.5B model, a direct wall-clock analysis shows that CROPI reaches 70% targeted accuracy with a 2.17x speedup over standard GRPO; we report the corresponding convergence and fixed-time comparisons in Appendix E.4.
@@
Table 2: Time costs of data selection and training for CROPI in each phase. We denote the number of prompts to process (gradient computation, projection, cosine similarity) in data selection and the optimization steps in the training stage in brackets. Time cost is measured on an 8-GPU (NVIDIA H100) machine.
@@ 6 Analysis
This section provides an in-depth analysis of two key computational components of the CROPI framework: random projection and data selection. We also provide an empirical analysis of off-policy gradient estimation in POPI in Appendix G.
@@ 6.1 Analysis on Sparse Random Projection
As mentioned on Section 3, we use sparse random projection for the projection of full-parameter gradients. Specifically, we randomly select a proportion of dimension of the gradient to perform random projection to avoid computation over the entire high-dimensional gradient. We define this proportion as sparse ratio. However, through our experiments, we found that this dropout actually improves the preservation of inner products between gradient features after random projection.
@@
Figure 4 presents the experimental results under the 1.5B Setting. We observe that when we directly apply random projection to the full gradient (i.e. sparse ratio equals 1) , the precision@10% of similarity ranking is only around 13%, comparable to random selection. However, when the sparse ratio is 0.1, the precision@10% is significantly higher, reaching nearly 80%. This is a counter-intuitive phenomenon: under sparse random projection, the ranking preservation is actually better with less information. We hypothesize that this may be related to the presence of numerical noise in the gradients: random projection can amplify such noise. Sparsity, while masking some information, filters out much of the numerical noise and achieves a better signal-to-noise ratio around a sparse ratio of 0.1. We include more analysis in Appendix H.
@@ 6.2 Analysis on data selected by CROPI
In Appendix I, we provide a detailed breakdown of data sources, knowledge categories, and the diversity of selected data. These results demonstrate that CROPI not only achieves strong performance but also offers a degree of interpretability.
@@ 7 Conclusion
Numerical error in gradient computation. As we utilize float16 format to conduct gradient-related calculations, numerical error cannot be ignored and might be amplified by random projection. We introduce a random dropout operation before random projection as a way to obtain better rank preservation. Additional details can be found in Appendix H.
@@ Appendix B Derivation of Off-Policy Gradient
However, for multi-step trajectory generation, the variance of the importance weight $\rho_{\theta}(\tau|s_{0})$ can become extremely large. This high variance arises due to the product of many likelihood ratios over the steps, leading to instability and poor sample efficiency in the gradient estimate. We therefore adopt a more widely used off-policy gradient estimator from (Schulman et al., 2015, 2017) that uses importance sampling at the token level:
@@ Appendix C Related Works
C.1 Data Selection for RLVR
Recently, how to automatically select high-quality data for RLVR has become a research topic of increasing interest. According to the time granularity of data selection, existing methods can be roughly divided into three categories: global-level, batch-level and phase-level. The first is the global-level granularity, which selects data based on the initial policy over the entire dataset Zhao et al. (2025); Wang et al. (2025). Although this approach is easy to implement, it cannot capture dynamic checkpoint information and often requires warm-up training. The second granularity is at the batch-level, where data selection is performed within each training batch. This method can better capture the dynamics of model training; however, due to the lack of long-term planning, it often introduces high variance that can lead to unstable training. Representative methods include DAPO (Yu et al., 2025) and ODF (Bae et al., 2025). The third category is a compromise, where data selection is performed every certain number of training steps, which we refer to as phase-level (or curriculum) data selection. This approach strikes a balance between capturing training dynamics and enabling long-term planning, but may still face challenges such as unstable training distributions. An example of this method is Xi et al. (2024).
As for the utility estimator, previous works use various ways to measure the utility of each data point, mainly focusing on heuristics around the difficulty and uncertainty of the training query $s_{0}$. In order to select prompts with proper difficulties, some researchers use correctness rate of various trajectories as the signal of the difficulty (Li et al., 2025; Yu et al., 2025; Bae et al., 2025; Sun et al., 2025). As an example, Bae et al. (2025) suggests filtering data points with pass rate around $0.5$ to construct the batch for GRPO training. Zhao et al. (2025) estimate prompt difficulty based on the model’s confidence (likelihood), and select moderately difficult prompts for reinforcement learning (RL) training. Another heuristic involves selecting data according to the uncertainty arising from data perturbation. For instance, Wang et al. (2025) train with prompts whose associated historical trajectories exhibit the highest reward variance.
@@ C.2 Influence Functions for LLM Data Selection
Influence function (Koh and Liang, 2017) is a gradient-based data attribution method derived from variantion analysis of objective function. Influence functions and their variants have been widely applied in the Pre-Training (PT) and Supervised Fine-Tuning (SFT) stages due to their strong theoretical guarantees and empirical effectiveness Logan Engstrom (2024); Wang et al. (2024); Gu et al. (2024); Grosse et al. (2023). Logan Engstrom (2024) proposed the datamodel framework leveraging an efficient influence function estimation Park et al. (2023), which is inspired by influence function, to select pretraining data. LESS (Xia et al., 2024) is the first to introduce first-order influence functions into the post-training of large language models (LLMs), thereby improving the training efficiency of supervised fine-tuning (SFT). Wang et al. (2024) further refine LESS by extending it to the batch-level; however, its application remains limited to the SFT stage. Nonetheless, leveraging influence function theory to guide data selection for RLVR remains a challenging problem. This is mainly due to the fact that, in RL settings, rewards and gradients for the data typically require rollouts to obtain, which is computationally expensive for LLMs. A concurrent work Hu et al. (2025) explores online data selection using influence functions in RLHF and demonstrates promising empirical results across multiple settings. However, their setting differs substantially from RLVR, and the target function used for influence estimation is defined for a different training objective.